Maple: A Quick Reference

Prepared by: Douglas Meade Department of Mathematics University of South Carolina

January 1995 (February 1997, updated for Release 4)

Symbols and Abbreviations

Symbol  Description  Example

:=               assignment  f := x^2/y^3;

;

terminate command; display result

                                                        int( x^2, x );
:

terminate command; hide result

                                                        int( x^2, x ):
..

                 specify a range or interval  plot( t*exp(-2*t), t=0..3 );
{ }

set delimiter

                                                        { y, x, y };
[ ]

list delimiter

                                                        [ y, x, y ];
"

refers to previous result (double quote) Int( exp(x^2), x=0..1 ):

                                                        " = value( " );

``

(also ?strings)

string delimiter (back quote)

Note: single quote (´ ) is different

                                                        TITLE := ` A Pretty Plot` ;

´´

(also ?uneval)

delayed evaluation (single quote) Note: back quote (` ) is different

                                                        x := ´ x´ ;

not needed in normal usage

->

mapping (procedure) definition f := (x,y) -> x^2*sin(x-y);

                                                        f(Pi/2,0);

Mathematical Operations, Functions, and Constants

Symbol  Description  Example

+, -, *, /, ^

add, subtract, multiply, divide, power

                                                        x^(-4)*3+x/Pi;
sin, cos, tan,
cot, sec, csc

trigonometric functions sin( x-Pi/5 )-sec( x^2 );

arcsin, arccos,
arctan, arccot,
arcsec, arccsc

inverse trigonometric functions arctan( 2*theta );

exp

exponential function

                                                        exp( 2*x );
ln

natural logarithm

                                                        ln( x*y/2 );
log10

common logarithm (base 10)

                                                        log10( 1000 );
abs

absolute value

                                                        abs( (-3)^5 );

sqrt             square root  sqrt( 24 );

!                factorial  k!;

=, <>, <, <=, >, >=

equations and inequalities

Note: E no longer exists; use exp(1)

                                                        diff( y(x), x ) + x*y(x) = F(x);
                                                        exp(Pi) > Pi^exp(1);
Pi, I

π, i (mathematical constants)

                                                        exp( Pi*I );
infinity

infinity

                                                        int( x^(-2), x=1..infinity );

\underline{NOTES}:

• The document is also available on the World Wide Web; the Universal Resource Locator is

    http://www.math.sc.edu/~meade/maple/maple-ref/

• Please send comments, corrections, and suggestions for improvements to meade@math.sc.edu.